OpenStudy (anonymous):

Compute the relative change in the function f(x)=x^2+1 from x=2 to x=5

OpenStudy (freckles):

I think it is \[\frac{f(5)-f(2)}{f(2)}\]

OpenStudy (gorv):

just put the value in function at x=2 f(2)=2^2+1 similarly for x=5

OpenStudy (gorv):

now we have to find relative change so divide it by initial value that is at x=2